# Weijia Shi1Jaechan Lee1Yangsibo Huang2 Sadhika Malladi2Jieyu Zhao3Ari Holtzman4Daogao Liu1 Luke Zettlemoyer1Noah A. Smith1Chiyuan Zhang5

### v1

@@ 1 Introduction
Training language models (LMs) often involves using vast amounts of text data, which may inadvertently contain private and copyrighted content (Carlini et al., 2021; Henderson et al., 2023; Min et al., 2023). In real-world applications, data owners may demand that their data be removed from a trained language model due to privacy or copyright concerns, as mandated for example by the General Data Protection Regulation (GDPR, European Parliament & Council of the European Union, ). Moreover, recent copyright lawsuits (DOE 1 v. GitHub, Inc., N.D. Cal. 2022; Tremblay v. OpenAI, Inc.,, 2023) emphasize the need for removing copyrighted data from the model.
@@ 3 The MUSE Evaluation Benchmark
Table 1: Comparison with a previous benchmark: Unlike the previous benchmark TOFU (Maini et al., 2024), which evaluates unlearning on synthetic Q&A datasets, MUSE tackles real-world unlearning challenges: unlearning real-world large-scale corpus (30$\times$ larger) while taking into account six desiderata that are important to both data owners and deployers. More related works are discussed in Appendix 6.
@@
Scale (##\## tokens in forget set) 0.8M for News, 1.1M for Books 0.15M
@@ 3.1 Evaluation Metrics
As discussed previously, it is desirable that the unlearned model does not leak membership information indicating that $\mathcal{D}_{\textrm{forget}}$ was part of $\mathcal{D}_{\textrm{train}}$. To determine if a given example was used during training, membership inference attack (MIA) exploits distributional differences in certain statistics (e.g., loss) between training (member) and non-training (non-member) data: if the loss on the example is low, then it was likely used for training. Unlearning typically increases the loss on the example, but there are two possible ways that unlearning can fail to prevent privacy leakage: (1) under-unlearning, when the loss is not made large enough; and (2) over-unlearning, when the loss is made abnormally large. To accurately measure the privacy leakage, we employ Min-KProb (Shi et al., 2024a) , a state-of-the-art MIA method for LMs based on the loss, and compute the standard AUC-ROC score (Murakonda et al., 2021; Ye et al., 2022a) of discriminating $\mathcal{D}_{\textrm{forget}}$ (members) and $\mathcal{D}_{\textrm{holdout}}$ (non-members). [^3] An MIA algorithm compares its score to a given threshold to classify a given datapoint as a member or non-member. The AUC-ROC is a single value that summarizes the overall performance of the MIA algorithm by measuring its ability to discriminate between members and non-members across all possible thresholds. By comparing the AUC score with that of the retrained model, we define [^4] Generally, $\textsf{AUC}(f_{\textrm{retrain}};\mathcal{D}_{\textrm{forget}},\mathcal{D}_{\textrm{holdout}})\approx 0.5$, though sometimes there are intrinsic distribution shifts between $\mathcal{D}_{\textrm{forget}}$ and $\mathcal{D}_{\textrm{holdout}}$ that may bias the baseline away from 0.5.
@@ 6 Related Work
Machine unlearning is a long-running, well-studied topic. Several studies have explored exact unlearning, aiming to make the unlearned model ($f_{\textrm{unlearn}}$) exactly identical to the reference model ($f_{\textrm{retrain}}$). As expected, this can only be accomplished in simple models like SVMs (Cauwenberghs & Poggio, 2000; Tveit et al., 2003; Romero et al., 2007; Karasuyama & Takeuchi, 2010) or naive Bayes models (Cao & Yang, 2015). Another approach is to ensure that the unlearned model $f_{\textrm{unlearn}}$ is probabilistically indistinguishable from $f_{\textrm{retrain}}$ (Ginart et al., 2019; Guo et al., 2020), and this view of certifiable unlearning is closely related to differential privacy (Dwork et al., 2006b, a). This rigorous definition of unlearning has inspired several theoretical works that characterize the feasibility of unlearning in convex and non-convex models, but those proposed algorithms are too computationally costly to operate on modern-day LMs (Izzo et al., 2021; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021; Gupta et al., 2021). Several more tractable unlearning algorithms have been proposed (Borkan et al., 2019; Ginart et al., 2019; Thudi et al., 2022; Chourasia & Shah, 2023) with broader applications such as image classification (Ginart et al., 2019; Golatkar et al., 2020a), text-to-image generation (Gandikota et al., 2023; Zhang et al., 2023; Fan et al., 2023), Federated Learning (Liu et al., 2020; Che et al., 2023; Halimi et al., 2022) and Recommender Systems (Li et al., 2024b).
@@
Machine unlearning has recently found its way into language model applications. In §4, we discuss some standard unlearning methods based on parameter optimization, like the Gradient Ascent and its variance. Other notable non-training-based unlearning methods include localization-informed unlearning (Meng et al., 2022; Wu et al., 2023; Wei et al., 2024), which involves identifying model units (e.g., layers, neurons) closely related to the unlearning data or tasks and then locally editing and modifying the units. In-context unlearning (Pawelczyk et al., 2023) offers another approach, treating the model as a black box and modifying its output results using external knowledge.
@@
Evaluating machine unlearning methods for language model applications is also critical. Most previous studies have focused this evaluation on specific tasks such as question answering or sentence completion. For example, Eldan & Russinovich (2023) experiment with unlearning to forget Harry Potter books and demonstrate the effectiveness of their methods by showing that familiarity scores, measured through completion-based, token-probability-based, and question-answering evaluations, significantly decline post-unlearning. Lynch et al. (2024) further suggest comparing unlearned models with perfectly retrained models. Their evaluation finds that while familiarity scores with the forget set may drop post-unlearning, they still remain higher than those of the retrained model. The closest work to ours is TOFU (Maini et al., 2024), a benchmark featuring 200 synthetic author profiles, each with 20 question-answer pairs, divided into forget and retain sets. However, TOFU is small-scale (0.15M tokens) and lacks separation of original text and knowledge, which conflates the evaluation of unlearning verbatim text and knowledge (C1 and C2 in MUSE). Additionally, current evaluations focus on limited aspects of data owner expectations and do not adequately reflect real-world deployment considerations, such as scalability and potential sequential unlearning requests. In contrast, MUSE formally defines different unlearning scopes and corresponding metrics, resulting in a systematic six-way evaluation featuring both data owners’ and deployers’ expectations. The evaluation uses a large-scale corpus of over 6 million tokens, separated into verbatim text and knowledge sets. We also note that some of our findings align with previous evaluations. For example, our observation that over- or under-unlearn can exacerbate privacy leakage (§5.2) is consistent with the recent work by Hayes et al. (2024).

### v2

@@ 1 Introduction
Training language models (LMs) often involves using vast amounts of text data, which may inadvertently contain private and copyrighted content (Carlini et al., 2021; Henderson et al., 2023; Min et al., 2023; He et al., 2024). In real-world applications, data owners may demand that their data be removed from a trained language model due to privacy or copyright concerns, as mandated for example by the General Data Protection Regulation (GDPR, European Parliament & Council of the European Union, ). Moreover, recent copyright lawsuits (DOE 1 v. GitHub, Inc., N.D. Cal. 2022; Tremblay v. OpenAI, Inc.,, 2023) emphasize the need for removing copyrighted data from the model.
@@ 3 The MUSE Evaluation Benchmark
Table 1: Comparison with a previous benchmark: Unlike the previous benchmark TOFU (Maini et al., 2024), which evaluates unlearning on synthetic Q&A datasets, MUSE tackles real-world unlearning challenges: unlearning real-world large-scale corpus (22$\times$ larger) while taking into account six desiderata that are important to both data owners and deployers. More related works are discussed in Appendix 6.
@@
Scale (##\## tokens in forget set) 0.8M for News, 3.3M for Books 0.15M
@@ 3.1 Evaluation Metrics
As discussed previously, it is desirable that the unlearned model does not leak membership information indicating that $\mathcal{D}_{\textrm{forget}}$ was part of $\mathcal{D}_{\textrm{train}}$. To determine if a given example was used during training, membership inference attack (MIA) exploits distributional differences in certain statistics (e.g., loss) between training (member) and non-training (non-member) data: if the loss on the example is low, then it was likely used for training. As shown in Figure 2, unlearning typically increases the loss on the example, but there are two possible ways that unlearning can fail to prevent privacy leakage: (1) under-unlearning, when the loss is not made large enough; and (2) over-unlearning, when the loss is made abnormally large. To accurately measure the privacy leakage, we employ Min-KProb (Shi et al., 2024a) , a state-of-the-art MIA method for LMs based on the loss, and compute the standard AUC-ROC score (Murakonda et al., 2021; Ye et al., 2022a) of discriminating $\mathcal{D}_{\textrm{forget}}$ (members) and $\mathcal{D}_{\textrm{holdout}}$ (non-members). [^3] An MIA algorithm compares its score to a given threshold to classify a given datapoint as a member or non-member. The AUC-ROC is a single value that summarizes the overall performance of the MIA algorithm by measuring its ability to discriminate between members and non-members across all possible thresholds. By comparing the AUC score with that of the retrained model, we define [^4] Generally, $\textsf{AUC}(f_{\textrm{retrain}};\mathcal{D}_{\textrm{forget}},\mathcal{D}_{\textrm{holdout}})\approx 0.5$, though sometimes there are intrinsic distribution shifts between $\mathcal{D}_{\textrm{forget}}$ and $\mathcal{D}_{\textrm{holdout}}$ that may bias the baseline away from 0.5.
@@ 6 Related Work
Machine unlearning is a long-running, well-studied topic. Several studies have explored exact unlearning, aiming to make the unlearned model ($f_{\textrm{unlearn}}$) exactly identical to the reference model ($f_{\textrm{retrain}}$). As expected, this can only be accomplished in simple models like SVMs (Cauwenberghs & Poggio, 2000; Tveit et al., 2003; Romero et al., 2007; Karasuyama & Takeuchi, 2010) or naive Bayes models (Cao & Yang, 2015). Another approach is to ensure that the unlearned model $f_{\textrm{unlearn}}$ is probabilistically indistinguishable from $f_{\textrm{retrain}}$ (Ginart et al., 2019; Guo et al., 2020), and this view of certifiable unlearning is closely related to differential privacy (Dwork et al., 2006b, a). This rigorous definition of unlearning has inspired several theoretical works that characterize the feasibility of unlearning in convex and non-convex models, but those proposed algorithms are too computationally costly to operate on modern-day LMs (Izzo et al., 2021; Neel et al., 2021; Ullah et al., 2021; Sekhari et al., 2021; Gupta et al., 2021). Several more tractable unlearning algorithms have been proposed (Borkan et al., 2019; Ginart et al., 2019; Thudi et al., 2022; Chourasia & Shah, 2023) with broader applications such as image classification (Ginart et al., 2019; Golatkar et al., 2020a), text-to-image generation (Gandikota et al., 2023; Zhang et al., 2023; Fan et al., 2023), Federated Learning (Liu et al., 2020; Che et al., 2023; Halimi et al., 2022; Huang et al., 2022) and Recommender Systems (Li et al., 2024b).
@@
Machine unlearning has recently found its way into language model applications. In §4, we discuss some standard unlearning methods based on parameter optimization, like the Gradient Ascent and its variance. Other notable non-training-based unlearning methods include localization-informed unlearning (Meng et al., 2022; Wu et al., 2023; Wei et al., 2024a), which involves identifying model units (e.g., layers, neurons) closely related to the unlearning data or tasks and then locally editing and modifying the units. In-context unlearning (Pawelczyk et al., 2023) offers another approach, treating the model as a black box and modifying its output results using external knowledge.
@@
Evaluating machine unlearning methods for language model applications is also critical. Most previous studies have focused this evaluation on specific tasks such as question answering or sentence completion. For example, Eldan & Russinovich (2023) experiment with unlearning to forget Harry Potter books and demonstrate the effectiveness of their methods by showing that familiarity scores, measured through completion-based, token-probability-based, and question-answering evaluations, significantly decline post-unlearning. Lynch et al. (2024) further suggest comparing unlearned models with perfectly retrained models. Their evaluation finds that while familiarity scores with the forget set may drop post-unlearning, they still remain higher than those of the retrained model. Wei et al. (2024b) evaluate the feasibility of using unlearning techniques to prevent language models from generating copyrighted content. The closest work to ours is TOFU (Maini et al., 2024), a benchmark featuring 200 synthetic author profiles, each with 20 question-answer pairs, divided into forget and retain sets. However, TOFU is relatively small-scale (0.15M tokens) and focuses on the evaluation of question answering. Additionally, current evaluations focus on limited aspects of data owner expectations and do not adequately reflect real-world deployment considerations, such as scalability and potential sequential unlearning requests. In contrast, MUSE formally defines different unlearning scopes and corresponding metrics, resulting in a systematic six-way evaluation featuring both data owners’ and deployers’ expectations. The evaluation uses a large-scale corpus of over 6 million tokens, separated into verbatim text and knowledge sets. We also note that some of our findings align with previous evaluations. For example, our observation that over- or under-unlearn can exacerbate privacy leakage (§5.2) is consistent with the recent work by Hayes et al. (2024). Our findings align with the the concurrent study by Shumailov et al. (2024) showing that unlearning gives a false sense of security as unlearned knowledge can resurface through in-context learning.